# Scattering wave packets of hadrons in gauge theories: Preparation on a quantum computer

### v2

@@ Abstract
preprint: UMD-PP-024-02 , IQuS@UW-21-071
I Introduction
Scattering processes have long served as a unique probe into the subatomic world, and continue to be the focus of modern research in nuclear and high-energy physics. They led to the discovery of fundamental constituents of matter, and enabled verification of the Standard Model of particle physics—a quantum mechanical and relativistic description of the strong and electroweak interactions in nature. Present-day and future particle colliders will continue to shed light on the structure of matter in form of hadrons and nuclei (e.g., at the Electron-Ion Collider Deshpande et al. (2005); Accardi et al. (2016); Abdul Khalek et al. (2022); Achenbach et al. (2023)). They will also probe exotic phases of matter (e.g., at the Relativistic Heavy-Ion Collider Csernai (1994); Bzdak et al. (2020); Achenbach et al. (2023); Lovato et al. (2022)), and will search for new particles and interactions beyond the Standard Model (e.g., at the Large Hadron Collider Narain et al. (2022); Bruning et al. (2012); Shiltsev and Zimmermann (2021)). A wealth of analytical and numerical methods have been developed over several decades to confront theoretical predictions based on gauge field theories of the Standard Model with experimental scattering data. These methods have reached unprecedented accuracy and precision in both perturbative and non-perturbative regimes of couplings Workman et al. (2022); Brock et al. (1995); Altarelli (1989); Dokshitzer (1991); Aoki et al. (2022). A highlight of such progress is the use of lattice-gauge-theory (LGT) methods to conduct first-principles studies of hadrons from the underlying theory of the strong force, quantum chromodynamics (QCD) Aoki et al. (2022); Davoudi et al. (2022a); Kronfeld et al. (2022).
The LGT method involves working with quantum fields placed on discrete and finite spacetime volumes with an imaginary time, so as to employ efficient Monte-Carlo techniques to compute observables. The Euclidean nature of computations, nonetheless, precludes direct access to scattering amplitudes. Indirect methods exist, starting from finite-volume formalisms developed in Refs. Luscher (1986, 1991), but they have been limited to low-energy and low inelasticity processes Briceno et al. (2018); Hansen and Sharpe (2019); Davoudi et al. (2021a). High-energy scattering of hadrons and nuclei are substantially more complex due to the composite nature of the colliding particles and a plethora of asymptotic final-state particles that are often produced. Beyond asymptotic scattering amplitudes, the evolution of matter as a function of time elapsed after the collision holds the key to yet-not-fully understood mechanisms of fragmentation, hadronization, and thermalization in particle colliders and in early universe Andersson et al. (1983); Webber (1999); Accardi et al. (2009); Albino (2010); Begel et al. (2022); Berges et al. (2021). Unfortunately, perturbative and non-perturbative tools, with the aid of classical computing, have had limited success in providing a full first-principles description of scattering processes to date.
@@ I Introduction
Alternatively, one can resort to Hamiltonian simulation, whose real-time nature is deemed favorable for simulating scattering processes from first principles. On classical computers, Tensor-Network methods have proven efficient in simulating gauge theories in the Hamiltonian formalism Bañuls et al. (2018); Bañuls and Cichy (2020); Meurice et al. (2022), including for scattering processes in simple models Pichler et al. (2016); Rigobello et al. (2021); Belyansky et al. (2023); Milsted et al. (2022); Van Damme et al. (2021). However, the exponential growth of the Hilbert space as a function of system size, and the accumulation of unbounded entanglement in high-energy processes, are likely to make classical Hamiltonian simulation of gauge theories of the Standard Model infeasible. This motivates exploring the potential of simulating these theories on quantum hardware Bañuls et al. (2020); Halimeh et al. (2023); Klco et al. (2022); Bauer et al. (2023a, b); Di Meglio et al. (2023). After mapping degrees of freedom of the system of interest to those of quantum platforms, time evolution can proceed in a digital or analog mode or a hybrid of both. The digital mode, which is the focus of this work, builds the unitary representing the Hamiltonian evolution out of a universal set of gates. The analog mode engineers a simulator Hamiltonian to mimic the target Hamiltonian, which is then evolved continuously. A hybrid mode combines features of both for more flexibility and efficiency. Digital  Martinez et al. (2016); Klco et al. (2018); Nguyen et al. (2022); Mueller et al. (2023); Chakraborty et al. (2022); de Jong et al. (2022); Lamm et al. (2019); Shaw et al. (2020); Klco et al. (2020); Atas et al. (2021); Kan and Nam (2021); Davoudi et al. (2022b); Ciavarella et al. (2021); Atas et al. (2023); Paulson et al. (2021); Haase et al. (2021); Kane et al. (2022); Cohen et al. (2021); Murairi et al. (2022); Davoudi et al. (2023); Farrell et al. (2023a); Charles et al. (2024); Mildenberger et al. (2022); Halimeh et al. (2022); Zhang et al. (2023); Kane et al. (2023); Hariprakash et al. (2023), analog Zhou et al. (2022); Yang et al. (2020); Mil et al. (2020); Banerjee et al. (2012); Zohar et al. (2013a, b); Hauke et al. (2013); Wiese (2013); Marcos et al. (2013); Zohar et al. (2016); Yang et al. (2016); Bender et al. (2018); Davoudi et al. (2020); Luo et al. (2020); Notarnicola et al. (2020); Surace et al. (2020); Surace and Lerose (2021); Kasper et al. (2020); Aidelsburger et al. (2021); Andrade et al. (2022); Surace et al. (2023), and hybrid Davoudi et al. (2021b); González-Cuadra et al. (2022); Zache et al. (2023); Popov et al. (2023); Meth et al. (2023) schemes have been developed and implemented in recent years for increasingly more complex gauge theories. Most implementations, nonetheless, concern dynamics after a quantum quench Mitra (2018). A quench process involves preparing the simulation in a simple initial state and abruptly changing the Hamiltonian to the Hamiltonian of interest in order to create non-equilibrium conditions. In order to simulate scattering processes, however, one needs to initialize the quantum simulation in more complex states, such as particle wave packets.
@@
Jordan, Lee, and Preskill (JLP) pioneered studies of scattering in quantum field theories on a quantum computer Jordan et al. (2012, 2014). Within an interacting scalar field theory, they laid down a systematic procedure for state preparation, time evolution, and measurement of wave packets. In particular, they considered the S-matrix formalism where the fields only interact in a scattering region, and are asymptotically free in the far past and the far future. Figure 1 summarizes the JLP protocol. Here, the interacting wave packets are obtained by evolving the free wave packets with a time-dependent Hamiltonian where the interactions are turned on adiabatically. Theoretically, the success of this adiabatic interpolation depends on relevant energy gaps Farhi et al. (2000), hence difficulties are expected when the adiabatic path crosses phase transitions. It also requires extra care to compensate for the broadening of the wave packet during the preparation time Jordan et al. (2012). In practice, the protocol becomes infeasible as the required adiabatic evolution time often exceeds the decoherence time of the present quantum hardware. Under the adiabatic framework, it is possible to achieve better resource scalings, e.g., by using more natural bases for the Hilbert space in certain limits, such as low particle numbers Barata et al. (2021), or by utilizing efficient ways to traverse the adiabatic path in the Hamiltonian’s parameter space Cohen and Oh (2023). Still, state preparation remains a bottleneck for general Hamiltonian simulation of scattering processes, particularly when the scattering particles are not excitations of the fundamental fields, and are rather composite (bound) states, as is the case in confining gauge theories.
@@
Various proposals have been put forth in recent years to eliminate the need for adiabatic state preparation in scattering protocols. For example, in 1+1-dimensional systems, one may extract the elastic-scattering phase shift using real-time evolution in the early and intermediate stages of the collision using measured time delay of the wave packets Gustafson et al. (2021). It is also possible to indirectly extract information about scattering by resorting to the relations between finite-volume spectrum of the interacting theory and scattering phases shifts Luscher (1986, 1991); Luu et al. (2010). Here, quantum computers replace classical computers to compute the spectra, e.g., using hybrid classical-quantum schemes such as variational methods Peruzzo et al. (2014); McClean et al. (2016); Tilly et al. (2022), see e.g., Ref. Sharma et al. (2023). As already mentioned, this limits the scope of such studies to asymptotic amplitudes and low energies, and does not benefit from the full power of quantum computers. Alternatively, the full state preparation in the JLP protocol can be implemented via a variational quantum algorithm Liu et al. (2022), assuming that suitable variational ansatze can be found for interacting excitations, which may in general be non-trivial. In another scheme, one may calculate scattering amplitudes directly from $n$-point correlation functions using the Lehmann–Symanzik–Zimmermann reduction formalism Li et al. (2023); Briceño et al. (2023, 2021), as well as particle decay rates with the use of optical theorem Ciavarella (2020), circumventing the need for wave-packet preparation. Nonetheless, creating wave-packet collisions that mimic actual scattering experiments allows probing the real-time dynamics of scattering, including entanglement generation and phases of matter, which are not directly accessible in asymptotic scattering amplitudes. Another alternative to circumvent adiabatic wave-packet preparation is to construct wave-packet operators directly in the interacting theory Turco et al. (2023), taking advantage of the Haag-Ruelle scattering theory Haag (1958, 1962) developed within axiomatic framework of quantum field theory Duncan (2012), see also Ref. Kreshchuk et al. (2023) for a similar approach. Nonetheless, extensions of these frameworks to theories with massless excitations, including gauge theories, are yet to be developed. Recently, an algorithm was developed and implemented on a quantum hardware which prepares wave packets of (anti)fermions in a fermionic lattice field theory in 1+1 dimensions (D) Chai et al. (2023). Fermionic excitations are generated from the interacting vacuum, which is itself prepared by a variational algorithm. Still, a practical algorithm for preparation of bound-state wave packets in gauge theories is desired.
@@
Figure 2: (a) A schematic comparison of the approach of this work and adiabatic state preparation. The blue and magenta paths represent the wave-packet creation in the JLP formalism. In order to arrive at the interacting wave packet, the adiabatic evolution is applied to the free wave-packet state created by $a_{\Psi}^{\dagger}$, which acts on the free ground state, |$\ket{\Omega}_{\rm free}$. The algorithm proposed in this work, the green path, directly builds the interacting wave-packet creation operator $b_{\Psi}^{\dagger}$ that acts on the interacting ground state, |$\ket{\Omega}$. |$\ket{\Psi}_{\rm free(int)}^{\rm in}$ denotes the incoming wave packet in the free (interacting) theory. $\Psi(k)$ denotes the wave-packet profile and $k$ is the momentum, see Sec. II for details. (b) A summary of various steps of the algorithm of this work, as detailed in Sec. III.1.
@@
Capitalizing on the idea of creating wave packets directly in the interacting theory, and inspired by a construction using tensor-network methods to study scattering in the lattice Schwinger Model Rigobello et al. (2021), we propose an efficient hybrid classical-quantum algorithm that creates wave packets of bound excitations in confining gauge theories in 1+1 D. This method, therefore, requires no adiabatic interpolation. Figure 2 provides a birds-eye view of our state-preparation algorithm compared to the JLP algorithm. Specifically, we use an efficient variational ansatz for the interacting ground state and verify its efficiency in the case of a $Z_{2}$ LGT with fermionic matter in 1+1 D. We further optimize a physically motivated ansatz for the meson creation operator out of such an interacting vacuum, which exhibits high fidelity in both perturbative and non-perturbative regimes of the couplings, as verified in the case of the $Z_{2}$ and $U(1)$ LGTs in 1+1 D. Our algorithm demands only a number of entangling gates that is polynomial in the system size, and uses a single ancilla qubit, benefiting from recently developed algorithms based on singular-value decomposition of operators Davoudi et al. (2022b). The quantum circuit that prepares a single wave packet is constructed for the case of the $Z_{2}$ LGT. For 6 fermionic sites (12+1 qubits), this circuit is executed on Quantinuum’s hardware, System Model H1, a quantum computer based on trapped-ion technology. Both the algorithmic and experimental fidelities are analyzed, and various sources of errors are discussed. We further discuss observables that can be measured efficiently in experiment to verify the accuracy of the generated wave packet. While (controlled) approximations are made to achieve shallower circuits, and a rather simple noise mitigation is applied based on symmetry considerations, high fidelities are still achieved in this small demonstration. Hence, quantum simulation of hadron-hadron collisions in lower-dimensional gauge theories may be within the reach of the current generation of quantum hardware.
@@
This paper is organized as follows. In Sec. II.1, we introduce the 1+1-dimensional LGTs coupled to staggered fermions. We then specify an interacting creation-operator ansatz in such theories in Sec. II.2, and demonstrate its validity through a numerical study in the case of $Z_{2}$ and $U(1)$ LGTs in Sec. II.3. The state-preparation algorithm and the circuit design are detailed in Sec. III.1, with a focus on the case of $Z_{2}$ LGT given its lower simulation cost. Section III.2 includes our results on the creation of wave packets with the use of both numerical simulators and a quantum computer. We end in Sec. IV with a summary and outlook. A number of appendices are provided to provide further details on the ansatz validity, circuit performance, and quantum-emulator comparisons. All data associated with numerical optimizations and circuit implementations are provided in Supplemental Material  [^1] The supplemental material can also be found at https://bit.ly/2402-00840.
II An ansatz for mesonic wave packets in gauge theories in 1+1 D
@@
II.1 Models: $Z_{2}$ and $U(1)$ LGTs coupled to fermions in 1+1 D
@@
Here, $\Gamma=\big{\{}0,a,\cdots,(N-1)a\big{\}}$ is the set of lattice-site coordinates. $a$ is the lattice spacing and $N$ denotes the number of staggered sites (and is hence even). $\xi^{\dagger}_{n}$ ($\xi_{n}$) stands for the fermionic creation (annihilation) operator at site $n$. In the staggered formulation, first developed in Refs. Kogut and Susskind (1975); Banks et al. (1976), fermions (antifermions) live on the even (odd) sites of the lattice while the links host the gauge bosons. $U_{n}$ and $E_{n}$ are non-commuting conjugate operators representing the gauge-link and the electric-field operators on the link emanating from site $n$. $m_{f}\geq 0$ is the fermion mass and $\epsilon$ is the strength of the electric-field Hamiltonian, expressed with the function $f(E_{n})$ for generality. For example, in the case of the $Z_{2}$ LGT, $f(E_{n})=E_{n}$ while in the $U(1)$ case, $f(E_{n})=E_{n}^{2}$. For the rest of this paper, we set $a=1$. The continuum limit is, therefore, realized in the limit of $m_{f},\epsilon\to 0$.
@@
Figure 3: The degrees of freedom in the $Z_{2}$ LGT are shown in (a). A lattice site $n$ (taken to be even in this figure) and its neighboring lattice site $n+1$ are denoted by circles, and the gauge link connecting them is denoted by a line. A fermion (antifermion) at $n$ $(n+1)$ is represented by an occupied (unoccupied) lattice site, which is shown here by $1$ $(0)$ in a green circle, and the absence of a fermion (antifermion) at $n$ $(n+1)$ is shown by 00 $(1)$ in a red circle. The corresponding fermion (antifermion) mass term in Eq. (1) is written above the lattice site $n$ $(n+1)$. The spin-$\frac{1}{2}$ hardcore boson residing on the link is indicated by a blue arrow along with its electric-field Hamiltonian term given in Eq. (1). For $\epsilon<0$ which is used in the simulations of this work and assumed in the figure, the down spin (green link) has higher electric-field energy than the up spin (red link). The action of the operator $\xi^{\dagger}_{n}\tilde{\sigma}^{\textbf{x}}_{n}\xi_{n+1}$, which is a part of the fermion hopping Hamiltonian in Eq. (1) for a $Z_{2}$ LGT, is depicted in (b). The upper (lower) dotted box in (c) shows configurations that satisfy the Gauss’s law given in Eq. (2) for the case of the $Z_{2}$ LGT at an even (odd) lattice site. Three different states for a $Z_{2}$ LGT with PBCs and 6 lattice sites are shown in (d). The top picture represents the strong-coupling vacuum state, |$\ket{\Omega}_{0}$, while the middle (bottom) picture depicts a forward-wrapped (backward-wrapped) 3-length bare-meson state created by the action of bare mesonic operators, defined in Sec. II.2, on |$\ket{\Omega}_{0}$.
@@
$G_{n}\ket{\psi_{\text{phys}}}=g\ket{\psi_{\text{phys}}}~{}\forall n,$ (2)
@@
II.2 Ansatz for interacting wave-packet creation operator

### v3

@@ Abstract
1 Introduction
Scattering processes have long served as a unique probe into the subatomic world, and continue to be the focus of modern research in nuclear and high-energy physics. They led to the discovery of fundamental constituents of matter, and enabled verification of the Standard Model of particle physics—a quantum mechanical and relativistic description of the strong and electroweak interactions in nature. Present-day and future particle colliders will continue to shed light on the structure of matter in form of hadrons and nuclei (e.g., at the Electron-Ion Collider [1, 2, 3, 4]). They will also probe exotic phases of matter (e.g., at the Relativistic Heavy-Ion Collider [5, 6, 4, 7]), and will search for new particles and interactions beyond the Standard Model (e.g., at the Large Hadron Collider [8, 9, 10]). A wealth of analytical and numerical methods have been developed over several decades to confront theoretical predictions based on gauge field theories of the Standard Model with experimental scattering data. These methods have reached unprecedented accuracy and precision in both perturbative and non-perturbative regimes of couplings [11, 12, 13, 14, 15]. A highlight of such progress is the use of lattice-gauge-theory (LGT) methods to conduct first-principles studies of hadrons from the underlying theory of the strong force, quantum chromodynamics (QCD) [15, 16, 17].
The LGT method involves working with quantum fields placed on discrete and finite spacetime volumes with an imaginary time, so as to employ efficient Monte-Carlo techniques to compute observables. The Euclidean nature of computations, nonetheless, precludes direct access to scattering amplitudes. Indirect methods exist, starting from finite-volume formalisms developed in Refs. [18, 19], but they have been limited to low-energy and low inelasticity processes [20, 21, 22]. High-energy scattering of hadrons and nuclei are substantially more complex due to the composite nature of the colliding particles and a plethora of asymptotic final-state particles that are often produced. Beyond asymptotic scattering amplitudes, the evolution of matter as a function of time elapsed after the collision holds the key to yet-not-fully understood mechanisms of fragmentation, hadronization, and thermalization in particle colliders and in early universe [23, 24, 25, 26, 27, 28]. Unfortunately, perturbative and non-perturbative tools, with the aid of classical computing, have had limited success in providing a full first-principles description of scattering processes to date.
@@ 1 Introduction
Alternatively, one can resort to Hamiltonian simulation, whose real-time nature is deemed favorable for simulating scattering processes from first principles. On classical computers, Tensor-Network methods have proven efficient in simulating gauge theories in the Hamiltonian formalism [29, 30, 31], including for scattering processes in simple models [32, 33, 34, 35, 36]. However, the exponential growth of the Hilbert space as a function of system size, and the accumulation of unbounded entanglement in high-energy processes, are likely to make classical Hamiltonian simulation of gauge theories of the Standard Model infeasible. This motivates exploring the potential of simulating these theories on quantum hardware [37, 38, 39, 40, 41, 42]. After mapping degrees of freedom of the system of interest to those of quantum platforms, time evolution can proceed in a digital or analog mode or a hybrid of both. The digital mode, which is the focus of this work, builds the unitary representing the Hamiltonian evolution out of a universal set of gates. The analog mode engineers a simulator Hamiltonian to mimic the target Hamiltonian, which is then evolved continuously. A hybrid mode combines features of both for more flexibility and efficiency. Digital  [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70], analog [71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91], and hybrid [92, 93, 94, 95, 96] schemes have been developed and implemented in recent years for increasingly more complex gauge theories. Most implementations, nonetheless, concern dynamics after a quantum quench [97]. A quench process involves preparing the simulation in a simple initial state and abruptly changing the Hamiltonian to the Hamiltonian of interest in order to create non-equilibrium conditions. In order to simulate scattering processes, however, one needs to initialize the quantum simulation in more complex states, such as particle wave packets.
@@
Jordan, Lee, and Preskill (JLP) pioneered studies of scattering in quantum field theories on a quantum computer [98, 99]. Within an interacting scalar field theory, they laid down a systematic procedure for state preparation, time evolution, and measurement of wave packets. In particular, they considered the S-matrix formalism where the fields only interact in a scattering region, and are asymptotically free in the far past and the far future. Figure 1 summarizes the JLP protocol. Here, the interacting wave packets are obtained by evolving the free wave packets with a time-dependent Hamiltonian where the interactions are turned on adiabatically. Theoretically, the success of this adiabatic interpolation depends on relevant energy gaps [100], hence difficulties are expected when the adiabatic path crosses phase transitions. It also requires extra care to compensate for the broadening of the wave packet during the preparation time [98]. In practice, the protocol becomes infeasible as the required adiabatic evolution time often exceeds the decoherence time of the present quantum hardware. Under the adiabatic framework, it is possible to achieve better resource scalings, e.g., by using more natural bases for the Hilbert space in certain limits, such as low particle numbers [101], or by utilizing efficient ways to traverse the adiabatic path in the Hamiltonian’s parameter space [102]. Still, state preparation remains a bottleneck for general Hamiltonian simulation of scattering processes, particularly when the scattering particles are not excitations of the fundamental fields, and are rather composite (bound) states, as is the case in confining gauge theories.
@@
Various proposals have been put forth in recent years to eliminate the need for adiabatic state preparation in scattering protocols. For example, in 1+1-dimensional systems, one may extract the elastic-scattering phase shift using real-time evolution in the early and intermediate stages of the collision using measured time delay of the wave packets [103]. It is also possible to indirectly extract information about scattering by resorting to the relations between finite-volume spectrum of the interacting theory and scattering phases shifts [18, 19, 104]. Here, quantum computers replace classical computers to compute the spectra, e.g., using hybrid classical-quantum schemes such as variational methods [105, 106, 107], see e.g., Ref. [108]. As already mentioned, this limits the scope of such studies to asymptotic amplitudes and low energies, and does not benefit from the full power of quantum computers. Alternatively, the full state preparation in the JLP protocol can be implemented via a variational quantum algorithm [109], assuming that suitable variational ansatze can be found for interacting excitations, which may in general be non-trivial. In another scheme, one may calculate scattering amplitudes directly from $n$-point correlation functions using the Lehmann–Symanzik–Zimmermann reduction formalism [110, 111, 112], as well as particle decay rates with the use of optical theorem [113], circumventing the need for wave-packet preparation. Nonetheless, creating wave-packet collisions that mimic actual scattering experiments allows probing the real-time dynamics of scattering, including entanglement generation and phases of matter, which are not directly accessible in asymptotic scattering amplitudes. Another alternative to circumvent adiabatic wave-packet preparation is to construct wave-packet operators directly in the interacting theory [114], taking advantage of the Haag-Ruelle scattering theory [115, 116] developed within axiomatic framework of quantum field theory [117], see also Ref. [118] for a similar approach. Nonetheless, extensions of these frameworks to theories with massless excitations, including gauge theories, are yet to be developed. Recently, an algorithm was developed and implemented on a quantum hardware which prepares wave packets of (anti)fermions in a fermionic lattice field theory in 1+1 dimensions (D) [119]. Fermionic excitations are generated from the interacting vacuum, which is itself prepared by a variational algorithm. Still, a practical algorithm for preparation of bound-state wave packets in gauge theories is desired.
@@
Figure 2: (a) A schematic comparison of the approach of this work and adiabatic state preparation. The blue and magenta paths represent the wave-packet creation in the JLP formalism. In order to arrive at the interacting wave packet, the adiabatic evolution is applied to the free wave-packet state created by $a_{\Psi}^{\dagger}$, which acts on the free ground state, |$\ket{\Omega}_{\rm free}$. The algorithm proposed in this work, the green path, directly builds the interacting wave-packet creation operator $b_{\Psi}^{\dagger}$ that acts on the interacting ground state, |$\ket{\Omega}$. |$\ket{\Psi}_{\rm free(int)}^{\rm in}$ denotes the incoming wave packet in the free (interacting) theory. $\Psi(k)$ denotes the wave-packet profile and $k$ is the momentum, see Sec. 2 for details. (b) A summary of various steps of the algorithm of this work, as detailed in Sec. 3.1.
@@
Capitalizing on the idea of creating wave packets directly in the interacting theory, and inspired by a construction using tensor-network methods to study scattering in the lattice Schwinger Model [33], we propose an efficient hybrid classical-quantum algorithm that creates wave packets of bound excitations in confining gauge theories in 1+1 D. This method, therefore, requires no adiabatic interpolation. Figure 2 provides a birds-eye view of our state-preparation algorithm compared to the JLP algorithm. Specifically, we use an efficient variational ansatz for the interacting ground state and verify its efficiency in the case of a $Z_{2}$ LGT with fermionic matter in 1+1 D. We further optimize a physically motivated ansatz for the meson creation operator out of such an interacting vacuum, which exhibits high fidelity in both perturbative and non-perturbative regimes of the couplings, as verified in the case of the $Z_{2}$ and $U(1)$ LGTs in 1+1 D. Our algorithm demands only a number of entangling gates that is polynomial in the system size, and uses a single ancilla qubit, benefiting from recently developed algorithms based on singular-value decomposition of operators [54]. The quantum circuit that prepares a single wave packet is constructed for the case of the $Z_{2}$ LGT. For 6 fermionic sites (12+1 qubits), this circuit is executed on Quantinuum’s hardware, System Model H1, a quantum computer based on trapped-ion technology. Both the algorithmic and experimental fidelities are analyzed, and various sources of errors are discussed. We further discuss observables that can be measured efficiently in experiment to verify the accuracy of the generated wave packet. While (controlled) approximations are made to achieve shallower circuits, and a rather simple noise mitigation is applied based on symmetry considerations, high fidelities are still achieved in this small demonstration. Hence, quantum simulation of hadron-hadron collisions in lower-dimensional gauge theories may be within the reach of the current generation of quantum hardware.
@@
This paper is organized as follows. In Sec. 2.1, we introduce the 1+1-dimensional LGTs coupled to staggered fermions. We then specify an interacting creation-operator ansatz in such theories in Sec. 2.2, and demonstrate its validity through a numerical study in the case of $Z_{2}$ LGT in Sec. 2.3. The state-preparation algorithm and the circuit design are detailed in Sec. 3.1, with a focus on the case of $Z_{2}$ LGT given its lower simulation cost. Section 3.2 includes our results on the creation of wave packets with the use of both numerical simulators and a quantum computer. We end in Sec. 4 with a summary and outlook. A number of appendices are provided to provide further details on the ansatz validity, circuit performance, and quantum-emulator comparisons. All data associated with numerical optimizations and circuit implementations are provided in Supplemental Material  [^1] The supplemental material can also be found at https://bit.ly/2402-00840..
2 An ansatz for mesonic wave packets in gauge theories in 1+1 D
@@
2.1 Models: $Z_{2}$ and $U(1)$ LGTs coupled to fermions in 1+1 D
@@
Here, $\Gamma=\big{\{}0,a,\cdots,(N-1)a\big{\}}$ is the set of lattice-site coordinates. $a$ is the lattice spacing and $N$ denotes the number of staggered sites (and is hence even). $\xi^{\dagger}_{n}$ ($\xi_{n}$) stands for the fermionic creation (annihilation) operator at site $n$. In the staggered formulation, first developed in Refs. [120, 121], fermions (antifermions) live on the even (odd) sites of the lattice while the links host the gauge bosons. $U_{n}$ and $E_{n}$ are non-commuting conjugate operators representing the gauge-link and the electric-field operators on the link emanating from site $n$. $m_{f}\geq 0$ is the fermion mass and $\epsilon$ is the strength of the electric-field Hamiltonian, expressed with the function $f(E_{n})$ for generality. For example, in the case of the $Z_{2}$ LGT, $f(E_{n})=E_{n}$ while in the $U(1)$ case, $f(E_{n})=E_{n}^{2}$. For the rest of this paper, we set $a=1$. The continuum limit is, therefore, realized in the limit of $m_{f},\epsilon\to 0$.
@@
Figure 3: The degrees of freedom in the $Z_{2}$ LGT are shown in (a). A lattice site $n$ (taken to be even in this figure) and its neighboring lattice site $n+1$ are denoted by circles, and the gauge link connecting them is denoted by a line. A fermion (antifermion) at $n$ $(n+1)$ is represented by an occupied (unoccupied) lattice site, which is shown here by $1$ $(0)$ in a green circle, and the absence of a fermion (antifermion) at $n$ $(n+1)$ is shown by 00 $(1)$ in a red circle. The corresponding fermion (antifermion) mass term in Eq. (1) is written above the lattice site $n$ $(n+1)$. The spin-$\frac{1}{2}$ hardcore boson residing on the link is indicated by a blue arrow along with its electric-field Hamiltonian term given in Eq. (1). For $\epsilon<0$ which is used in the simulations of this work and assumed in the figure, the down spin (green link) has higher electric-field energy than the up spin (red link). The action of the operator $\xi^{\dagger}_{n}\tilde{\sigma}^{\textbf{x}}_{n}\xi_{n+1}$, which is a part of the fermion hopping Hamiltonian in Eq. (1) for a $Z_{2}$ LGT, is depicted in (b). The upper (lower) dotted box in (c) shows configurations that satisfy the Gauss’s law given in Eq. (2) for the case of the $Z_{2}$ LGT at an even (odd) lattice site. Three different states for a $Z_{2}$ LGT with PBCs and 6 lattice sites are shown in (d). The top picture represents the strong-coupling vacuum state, |$\ket{\Omega}_{0}$, while the middle (bottom) picture depicts a forward-wrapped (backward-wrapped) 3-length bare-meson state created by the action of bare mesonic operators, defined in Sec. 2.2, on |$\ket{\Omega}_{0}$.
@@
$G_{n}\ket{\psi_{\text{phys}}}=g\ket{\psi_{\text{phys}}}\leavevmode\nobreak\ \forall n,$ (2)
@@
2.2 Ansatz for interacting wave-packet creation operator